# Structured Probabilistic Coding

### 21 Dec 2023 v1

@@ Introduction
Probabilistic embedding (Vilnis and McCallum 2015) learns a probability distribution that maps each data point to a distribution. The probabilistic approach can better capture the complexity and uncertainty of data, handle redundant information, and provide better discriminative representations. It has been applied to various domains such as computer vision (Oh et al. 2019; Shi and Jain 2019; Chang et al. 2020; Ramé and Cord 2021) and natural language processing (Mahabadi, Belinkov, and Henderson 2021; Wang et al. 2021; Hu et al. 2022a).
Most probabilistic embedding methods (Kingma and Welling 2014; Alemi et al. 2017; Higgins et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) are or can be built upon the principle of information bottleneck (IB) theory (Tishby, Pereira, and Bialek 1999; Tishby and Zaslavsky 2015). The principle aims to find a maximally compressed representation of the input that maximally preserves information about the output, striking a balance between compression and prediction. VIB (Alemi et al. 2017) is an efficient variational estimation method of the IB principle. For tractable application of IB in a continuous setting, Belghazi et al. (2018) propose a mutual information neural estimation method with IB principle, denoted as MINE-IB. And Ragonesi et al. (2021) employ MINE-IB to learn unbiased representations. Fischer (2020) and Ramé and Cord (2021) introduce a conditional mutual information term to alleviate the over- or under-compression issue of traditional IBs. Moreover, variational autoencoder (VAE) (Kingma and Welling 2014) is a special case of an unsupervised VIB and can be used to encourage disentangled representations (Higgins et al. 2017). Hu et al. (2022a) apply VAE to the masked pre-training process for learning diverse and well-formed contextual representations. Recently, An, Jammalamadaka, and Chong (2023) use the conditional entropy of the stochastic embedding as a confidence indicator and encourage the model to assign larger variance to more certain inputs.
@@ Conclusion
This paper proposes a new Structured Probabilistic Coding (SPC) framework to enhance the generalization ability of pre-trained language models for better language understanding. SPC adopts an encoder-only probabilistic coding technology to simultaneously perform information compression and task prediction. And a structured regularization is introduced to control probability distribution and promote class-level uniformity in the latent space. Experiments on 12 benchmarks show that SPC achieves the best performance on various classification and regression tasks. The results demonstrate that SPC can enhance pre-trained language models’ generalization capability, robustness to label noise, and the clustering quality of output representations.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (No. 2022YFC3302102). The authors thank the anonymous reviewers and the meta-reviewer for their helpful comments on the paper.
@@ Appendix A Experimental Setups
We conduct extensive experiments on various natural language understanding tasks including 10 classification tasks, and 2 regression tasks. The descriptions of each dataset and task are listed as follows:
@@
  • EmotionEval (Mohammad et al. 2018) involves predicting the most probable emoji of a tweet and is based on Emoji Prediction challenge conducted during SemEval-2018.
@@
  • SentiEval (Rosenthal et al. 2014) comes from SemEval 2017 and includes data from previous runs (2013, 2014, 2015, and 2016) of the same SemEval task. The goal is to determine if a tweet is positive, negative, or neutral.
@@
  • ISEAR (Scherer and Wallbott 1994) is from International Survey On Emotion Antecedents And Reactions project and contains reports on seven emotions each by close to 3000 respondents in 37 countries on all 5 continents. It aims to predict the emotion reaction.
  • MELD (Poria et al. 2019) contains multi-party conversation videos collected from Friends TV series, where two or more speakers are involved in a conversation. It is used to detect emotions in each utterance. [^5] The MELD dataset contains many types of context, including dialogue, speaker, and multi-modal data. Different from other task-oriented methods (Hu, Wei, and Huai 2021; Hu et al. 2022b), this work only considers the context-free textual utterance to better evaluate sentence classification performance.
  • GoEmotions (Demszky et al. 2020) is a corpus of comments from Reddit, with human annotations to 27 emotion categories or neutral. It is used for fine-grained emotion detection. [^6] The GoEmotions dataset contains nearly 16% multi-label data. In this paper, we remove this portion to better evaluate the multi-class classification performance.
Hyperparameter EmojiEval EmotionEval HatEval IronyEval OffensEval Trade-off weight $\beta$ 0.1 0.1 10 1 1 Trade-off weight $\gamma$ 10 0.1 0.1 10 0.1 Weight decay 0 0 0.001 0 0 Normalization False True False True False Hyperparameter SentiEval StanceEval ISEAR MELD GoEmotions Trade-off weight $\beta$ 0.01 1 0.1 1 0.001 Trade-off weight $\gamma$ 10 0.5 0.1 0.05 0.01 Weight decay 0 0 0.001 0 0 Normalization False False False True False
@@
Hyperparameter STS-B CLAIRE Trade-off weight $\beta$ 0.1 0.01 Weight decay 0 0 Normalization False False
@@
Figure 4: Performance against different trade-off weights $\beta$ of probabilistic coding for classification tasks. The experiments are conducted based on RoBERTa backbone. The grey line represents the results of CE baseline.
Figure 5: Performance of the optimal trade-off weight $\gamma$ for classification tasks. We experiment with RoBERTa backbone. The Y-axis refers to relative improvements between SPC and its variant SPC w/o Structure.
@@
  • STS-B (Cer et al. 2017) is a collection of sentence pairs drawn from news headlines, video and image captions, and natural language inference data. The semantic similarity task is to predict the semantic similarity score from 1 (very dissimilar) to 5 (very similar) given each pair.
  • CLAIRE (Roth, Anthonio, and Sauer 2022) datasets consists of manually clarified how-to guides from wikiHow [^7] https://www.wikihow.com/ with generated alternative clarifications and human plausibility judgements. The goal of plausible clarifications ranking task is to predict the continuous plausibility score on a scale from 1 (very implausible) to 5 (very plausible) given the clarification and its context.
@@
We report the detailed hyperparameter settings of SPC with RoBERTa backbone in Table 7 and Table 8. In the implementation of SPC, the hidden vector represents the output representation, and its dimension size is consistent with the dimension size of label space in each task. For each baseline, we fine-tune the key parameters following the original paper for fair comparison and to obtain corresponding optimal performance.
@@ Appendix B Supplementary Experiments
Performance against different weights of probabilistic coding
@@
Performance against optimal weight of task expression

### 23 Feb 2024 v4

@@ Introduction
Probabilistic embedding (Vilnis and McCallum 2015) learns a probability distribution that maps each data point to a distribution. The probabilistic approach can better capture the complexity and uncertainty of data, handle redundant information, and provide better discriminative representations. It has been applied to various domains such as computer vision (Oh et al. 2019; Shi and Jain 2019; Chang et al. 2020) and natural language processing (Mahabadi, Belinkov, and Henderson 2021; Wang et al. 2021; Hu et al. 2022).
Most probabilistic embedding methods (Kingma and Welling 2014; Alemi et al. 2017; Higgins et al. 2017; Fischer 2020; An, Jammalamadaka, and Chong 2023) are (or can be) built upon the principle of information bottleneck (IB) theory (Tishby, Pereira, and Bialek 1999; Tishby and Zaslavsky 2015). The principle aims to find a maximally compressed representation of the input that maximally preserves information about the output, striking a balance between compression and prediction. VIB (Alemi et al. 2017) is an efficient variational estimation method of the IB principle. For tractable application of IB in a continuous setting, Belghazi et al. (2018) propose a mutual information neural estimation method with IB principle, denoted as MINE-IB. And Ragonesi et al. (2021) employ MINE-IB to learn unbiased representations. Fischer (2020) and Ramé and Cord (2021) introduce a conditional mutual information term to alleviate the over- or under-compression issue of traditional IBs. Moreover, variational autoencoder (VAE) (Kingma and Welling 2014) is a special case of an unsupervised VIB and can be used to encourage disentangled representations (Higgins et al. 2017). Hu et al. (2022) apply VAE to the masked pre-training process for learning diverse and well-formed contextual representations. Recently, An, Jammalamadaka, and Chong (2023) use the conditional entropy of the stochastic embedding as a confidence indicator and encourage the model to assign larger variance to more certain inputs.
@@ Conclusion
This paper proposes a new supervised representation learning framework SPC, an encoder-only probabilistic coding technology with a structured regularization from the target space. By extracting compact and informative representations from input related to the target task, SPC can enhance the generalization ability of pre-trained language models for better language understanding. Firstly, an encoder-only probabilistic coding technology simultaneously performs variational encoding and task prediction. Then, a structured regularization is introduced to control probability distribution and promote uniformity across classes in the latent space. Experiments on 12 benchmarks show that SPC achieves the best performance on various classification and regression tasks. The results demonstrate that SPC can enhance the generalization capability, robustness to label noise, and the clustering quality of output representations.
@@ Appendix A Experimental Setups
We conduct extensive experiments on various natural language understanding tasks including 10 classification tasks, and 2 regression tasks, as shown in Table 6. The descriptions of each dataset and task are listed as follows:
@@
  • EmotionEval (Mohammad et al. 2018) involves detecting the emotion evoked by a tweet and is based on the Affects in Tweets conducted during SemEval-2018. Following Barbieri et al. (2020), the most common four emotions (i.e., anger, joy, sadness, and optimism) are selected as the label sets.
@@
  • SentiEval (Rosenthal, Farra, and Nakov 2017) comes from SemEval 2017 and includes data from previous runs (2013, 2014, 2015, and 2016) of the same task. The goal is to determine if a tweet is positive, negative, or neutral.
@@
  • ISEAR (Scherer and Wallbott 1994) is from International Survey On Emotion Antecedents And Reactions project and contains reports on seven emotions each by close to 3000 respondents in 37 countries on all 5 continents. It aims to predict the emotion reaction. Due to the lack of a predefined split in the original dataset, we randomly split the dataset into train/valid/test set in a ratio of 4:1:5 based on the label distribution.
  • MELD (Poria et al. 2019) contains multi-party conversation videos collected from Friends TV series, where two or more speakers are involved in a conversation. It is used to detect emotions in each utterance. [^5] The MELD dataset contains many types of context, including dialogue, speaker, and multi-modal data. Different from other task-oriented methods, e.g., DialogueCRN (Hu, Wei, and Huai 2021), this work only considers the context-free textual utterance to better evaluate sentence classification performance.
  • GoEmotions (Demszky et al. 2020) is a corpus of comments from Reddit, with human annotations to 27 emotion categories or neutral. It is used for fine-grained emotion detection. In this work, we remove all multi-label samples (nearly 16%) in the dataset to better evaluate the multi-class classification performance.
Dataset Task # Label # Train # Val # Test # Total
Classification
EmojiEval Emoji prediction 20 45,000 5,000 50,000 100,000
EmotionEval Social emotion detection 4 3,257 374 1,421 5,052
HatEval Hate speech detection 2 9,000 1,000 2,970 12,970
IronyEval Irony detection 2 2,862 955 784 4,601
OffensEval Offensive language detection 2 11,916 1,324 860 14,100
SentiEval Sentiment analysis 3 45,389 2,000 11,906 59.295
StanceEval Stance detection 3 2,620 294 1,249 4,163
ISEAR Emotion reaction prediction 7 3,066 767 3,833 7,666
MELD Conversational emotion recognition 7 9,989 1,109 2,610 13,708
GoEmotions Fine-grained emotion detection 28 36,308 4,548 4,591 45,447
Regression
STS-B Semantic similarity prediction - 7,000 1,500 1,400 9,900
CLAIRE Plausible clarification ranking - 19,975 2,500 2,500 24,975
Table 6: The statistics of all datasets.
Hyperparameter EmojiEval EmotionEval HatEval IronyEval OffensEval SentiEval StanceEval ISEAR MELD GoEmotions
Trade-off weight $\beta$ 0.1 0.1 10 1 1 0.01 1 0.1 1 0.001
Trade-off weight $\gamma$ 10 0.1 0.1 10 0.1 10 0.5 0.1 0.05 0.01
Weight decay 0 0 0.001 0 0 0 0 0.001 0 0
Layer normalization False True False True False False False False True False
@@
Hyperparameter STS-B CLAIRE
Trade-off weight $\beta$ 0.01 0.1
Weight decay 0 0
Layer normalization False False
@@
Figure 4: Performance against different trade-off weights $\beta$ of probabilistic coding for classification tasks. The experiments are conducted with RoBERTa backbone. The grey line represents the results of CE baseline.
Figure 5: Performance of the optimal trade-off weight $\gamma$ for classification tasks. We experiment with RoBERTa backbone. The Y-axis refers to relative improvements between SPC and its variant removing the structured regularization.
@@
  • STS-B (Cer et al. 2017) is a collection of English sentence pairs drawn from news headlines, video and image captions, and natural language inference data. The semantic similarity prediction task is to predict the semantic textual similarity score from 0 (very dissimilar) to 5 (very similar) given each sentence pair.
  • CLAIRE (Roth, Anthonio, and Sauer 2022) dataset consists of manually clarified how-to guides from wikiHow [^6] https://www.wikihow.com/ with generated alternative clarifications and human plausibility judgements. The goal of plausible clarifications ranking task is to predict the continuous plausibility score on a scale from 1 (very implausible) to 5 (very plausible) given the clarification and its context. In our experiments, a special token pair (i.e., <e>and </e>) is introduced as the boundary of filler words.
@@
We conduct experiments using an epoch number of $20$, a total batch size of $128$, and a maximum token length of $128$. The maximum patience for early stopping is set to 5 epochs. We report the detailed hyperparameter settings of SPC with RoBERTa backbone in Table 7 and Table 8. In the implementation of SPC, the hidden vector represents the output representation, and its dimension size is consistent with the dimension size of label space in each task.
For each baseline, we fine-tune the key parameters following the original paper for fair comparison and to obtain corresponding optimal performance. In addition, the temperature of GPT-3.5 is set to 0 for deterministic predictions.
@@ Appendix B Supplementary Experiments
Acknowledgements
This work was supported by the National Key Research and Development Program of China (No. 2022YFC3302102), and the National Funded Postdoctoral Researcher Program of China (No. GZC20232969). The authors thank the anonymous reviewers and the meta-reviewer for their helpful comments on the paper.